### v1

# Extracting Participation in Collective Action from Social Media

Abstract
Social media play a key role in mobilizing collective action, holding the potential for studying the pathways that lead individuals to actively engage in addressing global challenges. However, quantitative research in this area has been limited by the absence of granular and large-scale ground truth about the level of participation in collective action among individual social media users. To address this limitation, we present a novel suite of text classifiers designed to identify expressions of participation in collective action from social media posts, in a topic-agnostic fashion. Grounded in the theoretical framework of social movement mobilization, our classification captures participation and categorizes it into four levels: recognizing collective issues, engaging in calls-to-action, expressing intention of action, and reporting active involvement. We constructed a labeled training dataset of Reddit comments through crowdsourcing, which we used to train BERT classifiers and fine-tune Llama3 models. Our findings show that smaller language models can reliably detect expressions of participation (weighted F1=0.71), and rival larger models in capturing nuanced levels of participation. By applying our methodology to Reddit, we illustrate its effectiveness as a robust tool for characterizing online communities in innovative ways compared to topic modeling, stance detection, and keyword-based methods. Our framework contributes to Computational Social Science research by providing a new source of reliable annotations useful for investigating the social dynamics of collective action.
Introduction
The societal significance of online-mediated behavior change is growing, with grassroots, large-scale transformations being increasingly recognized as critical to addressing contemporary global challenges (Van Vugt 2009; Otto et al. 2020). While some online movements fail to yield tangible outcomes, in multiple occasions the Social Web has demonstrated its transformative potential in mobilizing collective action (Gerbaudo 2012) to achieve various political, societal, and economic goals (Yasseri et al. 2016; Chiovaro et al. 2021; Boulianne, Lalancette, and Ilkiw 2020; Lucchini et al. 2022).
Social movement theories suggest the presence of incremental pathways to collective action, where organized groups gradually mobilize individuals into becoming active participants (Van Stekelenburg and Klandermans 2013). Analyzing these trajectories—a critical focus for many Computational Social Science studies on opinion dynamics and behavior change—remains challenging due to the scarcity of reliable, large-scale ground truth data. The operationalization of social science theories relevant to collective participation, often overlooked in current literature on online behavior change (Pera, Morales, and Aiello 2023), could drive progress in this respect and support real-world applications.
To address these challenges, we combined insights from literature on mobilization and activism with modern Natural Language Processing (NLP) techniques to develop open-source text classifiers to detect expressions of participation in collective action from social media posts [^1] Data and code are available at https://tinyurl.com/4tj57wfk. Models are shared at https://huggingface.co/ariannap22.. Our classifiers identify the presence of participation and categorize it into one of four levels: (i) identifying collective problems and proposing solutions, (ii) engaging in calls-to-action, (iii) expressing intent to act, and (iv) reporting actual involvement in action. The tools are topic-agnostic, and therefore can be applied to large-scale platforms to generate extensive and granular ground truth data for various applications.
We evaluated four classification pipelines: a BERT-based classifier, a zero-shot Llama3 model, a Supervised Fine-Tuning (SFT) Llama3 model, and a Direct Preference Optimization (DPO) Llama3 model. We relied on Reddit posts manually annotated via crowdsourcing for supervised learning. While expressions of participation can be relatively rare in social media—especially for the highest participation levels—we found that modern data augmentation techniques can aid the creation of high-quality data sufficient for training.
Overall, we answer the following research questions:
RQ1. Can modern NLP reliably detect participation in collective action across topics?
RQ2. How can modern NLP models enhance existing methods and proxies to advance collective action analysis?
RQ3. Which Reddit communities exhibit higher levels of participation in collective action?
Our contributions include: (i) a framework to quantify participation levels in collective action, adaptable across topics and contexts; (ii) the operationalization of such framework with a human-annotated dataset and advanced NLP tools; (iii) validation experiments and a case study showcasing our method’s value compared to topic modeling, stance detection, and keyword-based methods tailored to capture topic-specific expressions of activism in language; and (iv) insights into how Reddit community demographics relate to collective action participation.
We found that small models can reliably quantify participation in collective action, often rivaling Large Language Models (LLMs). When applied to the binary classification task of identifying any expression of participation, the best BERT-based model achieved a macro F1 score of $0.71$, close to the $0.77$ of the best-performing LLM approach. In the multi-class classification of participation levels, the best BERT model obtained a macro F1 score of $0.44$, not far from the $0.52$ obtained by the best fine-tuned LLM. Notably, while fine-tuned LLMs excel at nuanced classifications, they demand significantly more computational resources, with inference times up to two orders of magnitude larger than what smaller transformer-based models require. The choice of model should therefore balance resource availability with task requirements.
Our proposed method differs from traditional topic modeling and stance detection, offering an effective tool for analyzing diverse data sources. Applied to climate-related discussions on Reddit, it detects abundant expressions of participation in collective action in communities that would not be recognized as action-oriented by widely-used keyword-based proxies. When extended to the broader Reddit platform, the method reveals higher levels of collective action participation in communities with audiences that are predominantly adult, female, and left-leaning. Such insights demonstrate the potential of our approach to considerably broaden the scope of studies of collective action in online settings.
Theoretical Framework
Our aim is to identify expressions of participation in collective action from natural language. To achieve this, we first need to define collective action and then identify the forms of participation that are relevant to the context of our study.
Collective Action
A well-established definition, rooted in psychology, describes collective action as any effort aimed at improving the conditions of an in-group, regardless of the number of participants or the outcome of the action (Wright, Taylor, and Moghaddam 1990). However, such a definition has some limitations. It overlooks the critical link between collective action and group presence, focuses solely on collective motives while excluding normative and self-interest as plausible ones, and does not consider actions by non-members designed to improve the status or treatment of disadvantaged out-groups (Wright 2009). These limitations make the definition unsuitable for the context of online discussions, where individual choices and motives often intersect with collective processes, and users frequently advocate for the rights of others outside their immediate in-group.
Given the limitations of the mainstream definition and its restricted applicability to online spaces, we propose an alternative definition of collective action that integrates individual perspectives and acknowledges the role of the out-group. We define collective action by first describing the nature of the collective action problem at its core. Specifically, we characterize a collective action problem as one that meets the following literature-informed criteria:
  • Human-generated. The issue arises from social, political, or economic factors and affects a group of people (or an individual as a representative symptom of a broader issue) who are required to make choices in an interdependent situation (Ostrom 2010).
  • Current. The problem is either ongoing or rooted in the past but with clear and significant impact on the present or future. Problems confined to the past, with no ongoing or future implications, simply do not necessitate current action for their resolution.
  • With opportunity for action. The issue can potentially be mitigated through collective efforts (e.g., rallies) or coordinated individual actions (e.g., recycling) (Van Zomeren 2013).
  • Involving shared responsibilities. Individuals, whether part of the in-group or out-group, can take actions that contribute to addressing the issue (Van Zomeren and Iyer 2009).
Based on this framework, we define collective action as any online or offline effort that can be taken to mitigate a collective action problem. Such mitigating actions can range from education and consciousness-raising, lobbying, negotiation, and petition signing to more disruptive activities such as boycotts, strikes and riots (Wright 2009).
Mobilization to Define Participation in Action
Mobilization is the process of allocating resources (such as time, money, or skills) to support a cause. It serves as a prerequisite for collective action. Existing theories on mobilization help categorize collective action stages, relating both to the individual agency to act and to the socio-psychological factors driving participation (Klandermans 1984). The literature on mobilization within social movements identifies the fundamental divide between non-action and action as a key conceptual distinction (Wright 2009; Klandermans and Stekelenburg 2014). Operationalizing this distinction is especially valuable for applications aimed at quantifying collective action online, without focusing on differentiating action participation stages.
Understanding the transition from non-action to action requires examining two key processes: consensus mobilization and action mobilization. Consensus mobilization refers to the organic convergence of views and meanings within social groups, driven by framing—namely a set of active processes that construct a shared reality (Klandermans 1988). There are three types of framing relevant to consensus mobilization (Benford and Snow 2000): (i) diagnostic framing, defining problems; (ii) prognostic framing, proposing solutions; and (iii) motivational framing, creating a rationale for action. Action mobilization, by contrast, focuses on promoting active participation (Gamson 1975). It unfolds through four steps (Klandermans and Oegema 1987): (i) ensuring individuals sympathize with the cause, (ii) informing them about upcoming events, (iii) encouraging a willingness to act and (iv) removing barriers to ensure involvement. Together, these two mobilization processes form a pathway for collective action, encompassing steps such as issue identification, solution proposals, calls to action, intention to act, and empowerment for active involvement.
Traditionally, mobilization mechanisms have been studied through the lens of social movement organizations as mobilizing agents (Snow et al. 1986). In this work, we adopt the distinction between “mobilization” and “participation” proposed by Klandermans and Oegema (1987) and shift the focus to the audience as the mobilized subject. This perspective applies the nuances of consensus and action mobilization to individuals, interpreting their steps toward participation as the grassroots counterpart to social movements’ mobilization efforts.
We therefore categorize expressions of participation in collective action into four classes:
  1. 1.
    Problem-solution: Identifying a collective action issue, assigning blame, or proposing solutions.
  2. 2.
    Call-to-action: Urging others to join or support a cause.
  3. 3.
    Intention: Expressing willingness or interest in taking action.
  4. 4.
    Execution: Reporting involvement in collective action initiatives.
These categories are well-suited for studying online discourse and generalize across various collective action initiatives by identifying levels of participation rather than specific activities tied to a particular cause.
Methodological Framework
We extracted and labeled comments from Reddit communities focused on activism and rights to create a training set for supervised learning. Comments were annotated on a multi-level scale: Problem-solution, Call-to-action, Intention, and Execution, which were collectively assigned the binary label Participation in collective action. Comments not fitting these categories were labeled as None. We then trained models in two stages: first, to perform binary classification of participation, and second, to classify the specific level of participation.
Data Curation
Data collection
We began our analysis by considering the top 40,000 subreddits ranked by all-time subscriber count, from which we identified a subset of 73 subreddits that contained the terms “activism,” “activist,” or “rights” in their titles or descriptions. Following a manual review, we selected 42 subreddits pertinent to our study’s scope, containing a total of 120 million comments posted from the creation of these subreddits through December 2023 (see Section A.1 in Appendix A for the exclusion criteria and the list of selected subreddits).
To refine our analysis towards comments more likely related to collective action, we adopted a word-matching approach. Specifically, we used an existing LIWC-like corpus comprising 47 terms associated with collective action (Smith, McGarty, and Thomas 2018). We retained only comments with a minimum of two term matches. We then split the comments into sentences and keep the sentence with the highest frequency of these terms, along with one preceding and one following sentence to preserve the discussion context. For each subreddit, we split this final dataset of comments into possible candidates for training and test, totaling 2.7 million and 150,000 comments respectively.
Crowdsourcing
We used Amazon Mechanical Turk (MTurk) to annotate comments. Workers were presented with one comment at a time and instructed to assign one label from Problem-solution, Call-to-action, Intention, Execution, or None. Label definitions and illustrative examples were provided (see Section A.2 in Appendix A). Workers were instructed to prioritize higher levels of participation when multiple labels applied (e.g., Execution over Intention).
We selected a random sample of 50 comments from each of the 42 subreddits, resulting in 2,100 samples. Workers were paid \$0.18 per annotation, corresponding to a hourly rate of \$7.20 considering 90 seconds per annotation.
To ensure annotations of high quality, we added 672 control samples (30% of real samples). These included 8 manually selected comments and 20 synthetically generated comments per participation label (i.e., excluding None) using Llama3 8B Instruct (Dubey et al. 2024) (cf. Appendix C for the prompt). Control samples had clear, known labels to evaluate annotator performance.
Overall, we employed five quality-control strategies. First, we presented the tasks as images rather than HTML text to prevent easy task automation. Second, we only recruited “master” workers who had completed at least 5,000 annotations on MTurk with a minimum 95% acceptance rate. Third, we discarded all annotations from workers failing more than 50% of control samples. Fourth, we removed samples with fewer than two annotators after filtering. Finally, we excluded samples without a clear majority vote. The final inter-annotator agreement measured with Krippendorff’s alpha is 0.86. The resulting annotated dataset consists of 369 samples: 100 labeled as None and 269 as expressing Participation in collective action, with 202 labeled as Problem-solution, 44 as Call-to-action, 14 as Intention, and 9 as Execution.
Data Augmentation
To address data sparsity in the annotated data, particularly in classes representing higher levels of participation, we applied two data augmentation techniques (Table 1).
The first approach involved generating synthetic samples using Llama3 8B Instruct (Møller et al. 2024). Excluding the majority class (Problem-solution), we generated 20 synthetic examples for each comment in the Call-to-action, Intention, Execution, and None classes. This dataset is referred to as Syn A. We also created a variation focusing solely on the minority classes Intention and Execution (Syn I/E). All generated samples were manually reviewed for accuracy. The generation prompt is provided in Appendix C.
The second method used Reddit’s structural features. Assuming similar comments in analogous contexts share the same label, we excluded comments with a similarity higher than 0.95–likely content generated by bots–and retrieved the top 5% most similar comments from the original discussion thread for each labeled comment. These were annotated with the same label as the original. Similarity was measured using Sentence-BERT (Reimers 2019) with a rolling window to handle truncation issues. The concatenation of the crowdsourced annotated dataset and this Reddit-informed extension is referred to as Ext.
Table 1: N. of participation labels by augmentation type.
Type P-Sol C-Act Intent Exec None
CS 202 44 14 9 100
CS + Syn I/E 202 44 275 182 100
CS + Syn A 202 234 275 182 240
Ext 553 175 24 16 285
Ext + Syn I/E 553 175 285 189 285
Test Set
To construct the test set, we randomly sampled 20 comments from each of the 42 selected subreddits, using the pool of 150,000 comments designated for testing. To ensure diversity, we limited the sampling to one comment per unique author-discussion thread combination. Consequently, some subreddits may have fewer than 20 comments in the test set. Consistent with the training set methodology, we extracted the sentence with the highest frequency of action-oriented terms (Smith, McGarty, and Thomas 2018), along with the preceding and following sentences. The final test set consists of 809 comments.
To ensure a high standard of annotation quality, we developed an annotation codebook (see Appendix A, Section A.3). This codebook further details the labeling guidelines established for the training set but it was deemed too complex for crowdworkers to use effectively within their time constraints. Both authors initially annotated 80 comments, achieving a Krippendorff’s alpha of 0.69. Discrepancies, mainly between Problem-solution and None, were resolved through discussion. The first author then annotated the remaining samples. In the end, 600 samples were labeled as None and 209 as expressing Participation in collective action, distributed as follows: 146 Problem-solution, 39 Call-to-action, 11 Intention, and 13 Execution.
Classification Models
We adopted a layered learning framework with a two-stage classification pipeline. First, a binary classification model determines whether a comment expresses participation in collective action. If participation is detected, a multi-class classification model identifies its level. This framework is designed to optimize the accuracy of the initial binary classification stage, which is critical for downstream tasks. This approach is common in multi-class classification with sparse labels (Cerqueira et al. 2023) and theoretically aligns with the collective action literature, emphasizing the distinction between participation and non-participation (Klandermans and Stekelenburg 2014) and accommodating potential use cases that may only necessitate a binary classification.
We evaluated four state-of-the-art Natural Language Processing (NLP) pipelines (Møller and Aiello 2024): (i) a BERT-based classifier fine-tuned on Reddit, (ii) a zero-shot Large Language Model (LLM) classifier, (iii) a Supervised Fine-Tuned LLM (SFT), and (iv) an LLM using Direct Preference Optimization (DPO). We also included word matching and centroid classification as baseline models. Modeling details on training and prompt definitions are detailed in Section A.4 of Appendix A and in Appendix C.
Dictionary (Dict)
For binary classification, we used the dictionary of collective action words (Smith, McGarty, and Thomas 2018) to calculate the fraction of such words in the text. We then selected the threshold maximizing the difference between the True Positive Rate (TPR) and False Positive Rate (FPR) on the training set.
Centroid classifier (Centr)
For both the binary and the multi-class tasks, we considered a centroid-based method, which has been used for similar classification tasks due to its simplicity and fast inference (Park, Hong, and Kim 2020). This method computes the semantic similarity between a sample and the embedding centroid of text samples of each specific class, and selects the class with the closest centroid. We implemented it using the cosine similarity between the Sentence BERT embeddings of text samples (Reimers 2019).
BERT Classifier (BERT)
We fine-tuned a RoBERTa model (Liu 2019) on the full set of 2.7M Reddit comments reserved for training, with the aim of adapting the language model to recognizing linguistic patterns in Reddit comments related to collective action. The model is then trained for both binary and multi-class classification tasks by considering a weighted loss function due to class imbalance.
Zero-shot LLM (ZS)
In a zero-shot setting, a LLM classifies text based on its internal understanding of the task. We designed the prompt using AI-enhanced knowledge prompting (Liu et al. 2021), leveraging Llama3 8B Instruct for label definitions and GPT-4 for prompt refinement. The classification of participation in collective action inherently relies on the definition of collective action and, by extension, the definition of a collective action problem. To explore this dependency, we experimented with prompts that either excluded or included a definition of the collective action problem (referred to as no def and def, respectively).
Fine-tuned Llama Models
We applied Supervised Fine-Tuning (SFT) to Llama3 8B Instruct by providing the model with a series of prompts that include: (i) fixed classification instructions specific to the task, along with all the allowed labels, and (ii) a set of labeled texts, derived from the crowdworkers’ annotated dataset. We followed current best practices by adopting Quantized Low-Rank Adaptation (QLoRA), an efficient fine-tuning technique (Dettmers et al. 2024). In this approach, the main model is frozen and quantized into a 4-bit representation, while separate low-rank matrices of gradients are learned during the fine-tuning process. These matrices are then combined with the frozen model during inference, weighted by a factor $\alpha$.
As an additional optimization step, we employed Direct Preference Optimization (DPO), a technique that updates the model’s weights based on explicit user preferences for one training example over another (Rafailov et al. 2024). During DPO, the model receives a prompt along with pairs of responses ranked by preference. The model’s weights are updated using cross-entropy loss to maximize the probability of generating the preferred example.
As with the zero-shot model, we tested prompts with and without a collective action problem definition (no def and def).
Results
Binary Task
Table 2 reports the performance of the binary classification task–the first step in our layered approach–measured in terms of F1 scores. The best-performing models were the zero-shot Llama3 model (both with and without the collective action problem definition in the prompt) and the RoBERTa classifier trained on the synthetically augmented training set. Across methods, the synthetic augmentation improved the classification performance considerably, up to $80\%$ for the macro F1 for RoBERTa. Fine-tuned LLMs underperformed simpler approaches.
The computational time required for training and inference in LLM fine-tuning was significantly higher compared to RoBERTa. Using a Tesla V100 GPU, the SFT training took 1.5 to 4 hours to complete depending on the selected augmented training set, the DPO training required 12.5 to 24 hours and the fine-tuned RoBERTa training only 4 minutes. Using a GTX 1080ti GPU, the average inference time per entry was 0.67s for the zero-shot approach, 1.09s for the fine-tuned LLM models and 0.007s for the RoBERTa classifier. Given the comparably good performance, the sensitivity of LLMs to prompt changes, which can lead to inconsistencies (Liu et al. 2023), the higher computational costs during inference and the complexity of hosting the models locally, we decided to select the RoBERTa model as the starting point for the multi-class step of the layered approach.
Table 2: F1 scores for the binary classification task. The three best-performing models are in bold.
Method Train. set Part. None Macro Avg. Weight. Avg.
Dict. 0.39 0.51 0.45 0.48
Centr. 0.49 0.63 0.56 0.60
BERT CS 0.44 0.28 0.36 0.32
Syn A 0.51 0.78 0.65 0.71
ZS no def 0.56 0.84 0.70 0.76
def 0.59 0.83 0.71 0.77
SFT CS (no def) 0.47 0.50 0.49 0.49
CS (def) 0.48 0.47 0.47 0.47
Syn A (no def) 0.50 0.59 0.54 0.57
Syn A (def) 0.51 0.59 0.55 0.57
DPO Syn A (def) 0.47 0.42 0.44 0.43
Table 3: Performance metrics for the multi-class classification task as a second step of the layered approach. The three best-performing models are in bold.
Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
Centr. 0.38 0.15 0.04 0.11 0.78 0.29 0.66
BERT CS 0.36 0.56 0.11 0.25 0.78 0.41 0.68
Syn A 0.36 0.62 0.17 0.29 0.78 0.44 0.68
ZS no def 0.34 0.54 0.09 0.46 0.78 0.44 0.67
def 0.38 0.39 0.09 0.36 0.78 0.40 0.67
SFT (def) CS 0.37 0.56 0.13 0.43 0.78 0.45 0.68
Ext 0.35 0.38 0 0 0.78 0.3 0.66
Ext + Syn I/E 0.36 0.39 0.2 0.3 0.78 0.41 0.67
CS + Syn I/E 0.36 0.55 0.17 0.42 0.78 0.46 0.68
CS + Syn A 0.37 0.64 0.29 0.5 0.78 0.52 0.69
DPO (def) CS 0.28 0.17 0 0.29 0.78 0.30 0.64
Ext 0.34 0.44 0 0 0.78 0.31 0.66
Ext + Syn I/E 0.34 0.26 0 0.12 0.78 0.3 0.65
CS + Syn I/E 0.37 0.62 0.21 0.52 0.78 0.5 0.69
CS + Syn A 0.38 0.58 0.29 0.5 0.78 0.51 0.69
Multi-class Task
In the second step, we trained classification models to identify the four levels of participation in collective action. During testing, we focused on comments selected in the first classification stage by the BERT-based classifier trained on synthetically augmented data.
Table 3 reports performances for the multi-class task, considering the inclusion of the collective action problem definition in the LLM prompts. The SFT Llama3 model, trained on crowdworker-annotated data with synthetic augmentation for all classes except the majority class (CS+Syn A), achieved the best overall performance. The DPO Llama3 model, also trained on annotated data with the two synthetic augmentation strategies, followed closely. Given that prompt variations and the limited number of Intention and Execution labels in the test set affect reliability for these classes, the RoBERTa classifier, trained on synthetically augmented data, remains a robust choice for this stage. However, the SFT and DPO models consistently outperformed others, particularly in classifying the minority classes Intention and Execution.
The ranking of models remained largely consistent whether or not the collective action problem definition was included in the LLM prompts or when models were applied directly to the multi-class task without a layered approach (see Appendix B, Tables B2 and B3).
Validation
We conducted four experiments to validate our proposed method. In particular, we focused on the first step of the layered pipeline, which is crucial to the overall classification task, and consider the BERT-based classifier trained on synthetically augmented data as reference. First, we compared our method to topic modeling to show that expressions of participation in collective action go beyond mere thematic categorization. Second, we contrasted our approach with stance detection, a task focused on identifying a position on a specific issue. Our method differs in its applicability across issues and in the fact that it focuses on behaviors rather than on opinions. Third, as our training sample selection was based on the collective action words dictionary (Smith, McGarty, and Thomas 2018), we examined the impact of this choice on the final classification results. Finally, we evaluated the binary classification of participation in collective action on a dataset that is distinct from Reddit.
(a) Datasets
(b) Topics
(c) Commitment
(d) Test Set Topics
(e) Stance
(f) Keywords
Figure 1: Validation of the proposed approach. (a-d) comparison with topic modeling; (e) comparison with stance detection; (f) impact on classification performance given by the use of collective action keywords to define training and test sets.
Topic Modeling
Our approach is (i) topic-agnostic, not distinguishing between different collective action topics, and (ii) capable of detecting participation in collective action even in subreddits not included in the training set. To demonstrate this, we analyzed three Reddit threads relevant to collective action: one from r/intj on COVID-19 vaccine intake (“Did you get the COVID vaccine? Why or why not?”, 679 comments), one from r/wallstreetbets on class action lawsuits in the GameStop short squeeze (“Robinhood App Hit With Class Action Suit After Trying to Shut Down GameStop Uprising”, 200 comments), and one from r/IronFrontUSA–part of our training set–on the importance of voting (“Posting this again to remind people that if you refuse to vote […] you share a huge amount of responsibility for what just happened”, 180 comments).
We collected all comments from these threads and applied BERTopic (Grootendorst 2022) to a shuffled concatenation of them. We performed a minor text pre-processing by removing English stopwords and the 20 most frequent words in the whole collection. This resulted in nine distinct topics (plus noise), of which the largest ones are: covid and its effects (332 comments), voting and democrats (165 comments), science and intjs (141 comments), robinhood, money, and GameStop (105 comments) (cf. Figure 1(b)). The remaining five topics are significantly smaller in size (with the biggest containing 31 comments) (cf. Table B4 and Figure B2 in Appendix B).
If our method were not topic-agnostic, Figure 1(c) would show a distribution similar to Figure 1(a). However, comments expressing participation in collective action do not follow a clear dataset dependency, although participation is featured less prominently in the r/wallstreetbets thread.
If our method’s detection capability were limited to its training set, no comments expressing participation would be detected in r/intj or r/wallstreetbets. However, at least some comments expressing participation were identified in all analyzed subreddits (see Figure 1(c)).
Additionally, if our method resembled topic modeling, we would expect topics extracted from the reference test set using BERTopic to be predominantly classified as either full participation in collective action or None. However, among the 14 identified topics, most are nearly evenly split between participation and None (see Figure 1(d)).
Stance Detection
In addition to being applicable across various topics, we demonstrate that our approach differs from stance detection, particularly in how stance classes relate to participation in collective action. To test this, we used a stance detection dataset related to global warming (Luo, Card, and Jurafsky 2020). This dataset includes sentences from news articles annotated for their stance towards the statement “Climate change/global warming is a serious concern,” with labels agree, disagree, and neutral.
We applied our binary classification approach to the dataset and analyzed the distribution of participation labels across stance classes. If our method were akin to stance detection, we would expect each stance class to align predominantly with either full participation or None. However, as shown in Figure 1(e), while there is some correlation between stance and participation, participation is detected in all stance classes at varying levels, indicating that stance detection cannot fully replace our method.
Impact of Collective Action Terms
We used the collective action words dictionary (Smith, McGarty, and Thomas 2018) to filter comments in both the training and test sets, as well as to extract relevant portions of each comment for classification models and annotators. To ensure our classifier does not rely solely on the presence of these words, we tested the classifier on four variations of the test set: (i) removal of action dictionary words, (ii) substitution of action dictionary words with random tokens from the model vocabulary, (iii) removal of random words, and (iv) substitution of random words with random tokens.
If collective action words influenced the classifier, we would expect different results for configurations (i) and (ii) compared to the original test set. However, as shown in Figure 1(f), the distribution of participation does not significantly differ under any of the tested conditions.
Classifying Parliamentary Debates
To demonstrate the broader applicability of our framework, we applied the classifier to detect participation in collective action in a subset of UK parliamentary debate transcripts (MySociety 2024). We selected debates from March 22nd, 25th, and 26th, 2024, totaling 1,000 text samples. Our approach effectively distinguishes messages expressing participation from those that do not, either by being unrelated to collective action or not identifying the issues as problematic (Table 4).
Table 4: Selected examples of parliamentary debates transcripts from the top and bottom 10 probabilities of class Participation in collective action.
Text
High probability 1. There is a desperate need for increased humanitarian support to Gaza. The UK, including the Ministry of Defence, is working collectively with allies […]
2. The Israeli Government have said that they want to “flood” Gaza with aid. Will my right hon. Friend assure the House that we will work with our partners globally to get more aid […]
3. […] the Prime Minister and I have been very proactive in speaking to and making multiple visits to the region. […] There is now a large-scale programme of using a pier to get food in […]
Low probability 1. I said that we are not going to be partisan in this debate, and the shadow Minister started in that vein, but my right hon. […]
2. The hon. Gentleman’s point is made more potent by the fact that the matters the ISC considers are not typically—in fact, not at all—partisan. […]
3. How many prisoners have been released early under the end of custody supervised licence scheme since October 2023.
Figure 2: Comparison of climate change-focused comment percentages and participation in collective action levels across subreddits, with x-axis values normalized using min-max scaling.
Climate Action
We now demonstrate the effectiveness of our method for detecting comments indicative of participation in collective action, comparing it to widely used approaches in the literature. We used climate action as a case study due to the extensive research in this area, particularly within Reddit (Oswald and Bright 2022; Treen et al. 2022). Previous studies have largely relied on subreddit selection and keyword filtering, assuming that the presence of certain keywords or subreddit participation directly reflects participation in climate action (Parsa et al. 2022; Lenti et al. 2024). In contrast, we show that subreddits selected based on keyword counts may not always align with actual participation in collective action.
We began with a potential set of the top 2,000 all-time subreddits ranked by subscriber count. Using a list of climate change-related substrings from (Parsa et al. 2022) and focusing on 2018, we filtered this set to include only subreddits containing climate-related comments, resulting in 1,597 subreddits. To further enhance our dataset, we added additional subreddits identified as relevant by Parsa et al. (2022) that were not part of the initial set, bringing the total to 1,609 subreddits. From these communities, we collected 719,667 keyword-matching comments posted in 2018, which were then analyzed with our best BERT-based binary classifier to detect those expressing participation in collective action. To streamline our analysis, we focused on 440 subreddits with at least 100 climate change-related comments during the specified timeframe. These subreddits were ranked based on two criteria: (i) the fraction of comments containing climate-related keywords and (ii) the fraction expressing a participation in collective action alongside climate change-related keywords. The rankings showed a moderate positive but weak correlation (Spearman $R=0.34$), with notable divergence at the top positions. Figure 2 highlights 24 subreddits where at least 40% of comments express participation in action, ranked by ascending percentage of such a measure. Subreddits like r/climate and r/environment, while highly ranked by climate-related mentions, did not appear in the top 10 for participation in collective action. Conversely, subreddits such as r/marchagainsttrump, r/bluemidterm2018, and r/democrats ranked very low for climate mentions but show a relatively high frequency of participation in collective action. Depending on the specific type of participation in collective action, detected through Llama3 fine-tuned on synthetically augmented data, different subreddits may be more or less suitable as data sources. For instance, r/ClimateOffensive, r/self, r/vegetarian and r/anticonsumption were the highest ranked in terms of Execution among the 24 visualized communities. This case study demonstrates how our framework identifies online communities relevant to climate activism more effectively than coarse metrics like subreddit popularity or keyword frequency.
Sociodemographic Dimensions of Participation
With the rise of online media in public discourse (Boulianne and Theocharis 2020), research increasingly seeks to characterize the socio-political attributes of social media users through embedding-based methods that distill semantic information from user content to infer attributes like age, income, and partisanship (Aceves and Evans 2024). These embeddings have been used to study political engagement by correlating key dimensions with outcomes such as collective decision-making, online mobilization, and advocacy for global causes (Israeli, Kremiansky, and Tsur 2022; Kumar et al. 2018; Lenti et al. 2024). However, such studies often rely on coarse proxies, such as subreddit participation as a stand-in for activism on Reddit (Lenti et al. 2024).
Our classification pipeline presents a new opportunity to examine whether the socio-political dimensions estimated by state-of-the-art embedding methods correlate with direct expressions of participation in collective action, and whether these correlations align with theoretical expectations. Specifically, we tested three hypothesis: (i) younger generations contribute significantly to activism, though their involvement may differ from older generations (Earl, Maher, and Elliott 2017), (ii) women and men tend to engage differently in political activism (Coffé and Bolzendahl 2010), and (iii) awareness of wealth inequality can drive people toward activism, with liberals more likely to engage in activism aimed at reducing inequalities (Hoyt et al. 2018).
To test these hypotheses, we used Reddit community embeddings (Waller and Anderson 2019) to characterize the socio-political position of subreddits. For the top 1,000 subreddits by all-time subscriber count, we combined these embeddings with the social dimensions defined by Waller and Anderson (2021) to compute scores estimating communities’ standing along three axes: age (young/old), gender (male/female), and partisanship (left-wing/right-wing). These scores were then compared to the fraction of comments expressing explicit participation in collective action, predicted through our BERT-based classifier trained on synthetically augmented data. The embeddings were derived from comments posted between 2005 and 2018 and, to meet computational constraints, the collective action analysis used a random 5% sample of comments published in May 2018, totaling 1.4 million.
While variability across subreddits is substantial, Figure 3 reveals three notable patterns. First, subreddits aligned more with the female pole show a higher fraction of comments expressing participation in action, supporting the hypothesis that gender is a dimension that partly explains activism patterns. Second, subreddits skewed towards older demographics also exhibit a higher fraction of such comments, suggesting meaningful contributions from older generations. Note that “older” here reflects platform-relative demographics and does not contradict the hypothesis that younger generations are highly active, considering that 64% of Reddit users are under 30 years old (Barthel et al. 2016). Last, there is a weaker, inverse relationship between conservative-leaning subreddits and the prevalence of participation-related comments, consistent with the hypothesis that liberals are more likely to engage in activism.
Figure 3: Fraction of Reddit comments showing participation in collective action across subreddits with varying socio-political tendencies. The socio-political scores are divided into equally sized bins based on sample quantiles.
Discussion and Conclusion
Drawing inspiration from the literature on collective action and mobilization, we contribute to the Computational Social Science community with new open-source text classifiers designed to accurately extract expressions of participation in collective action across topics. Our classification pipeline can operate in a layered fashion to capture different levels of participation. We achieve this by employing state-of-the-art Natural Language Processing models through a supervised learning approach, complemented by modern data augmentation techniques. We show that relatively high classification accuracy is attainable even with a limited number of manually-labeled examples. In our analysis of 61 different combinations of models and data augmentation strategies, we discovered that transformer-based small language models can compete effectively with more modern Large Language Models. Notably, our BERT-based classifier offers an optimal balance between performance, computational efficiency, and cost-effectiveness across both layers of classification (RQ1).
In applying our tool to data derived from Reddit discussions, we observed that our classification yields results that are orthogonal to those produced by traditional topic modeling and stance detection algorithms. When applied to climate-related communities, our tool reveals a profile of participation in collective action that differs significantly from the results obtained through conventional keyword matching methods. This highlights the potential of our approach to improve the text analytics pipeline over topic-specific approaches and coarser proxies of collective action previously employed in the literature (RQ2). Furthermore, our analysis indicates that discussions rich in expressions of participation in collective action are prevalent in forums characterized by more female, adult and left-leaning participants, in line with theoretical expectations (RQ3).
Our proposed method has some limitations. First, the volume of the annotated training set is relatively small. Although we have implemented data augmentation techniques to address this limitation, the collection of additional annotations could further enhance classification quality. Second, while the training set consists of crowdsourced annotations, our test set is annotated by domain experts. Although we prioritized the accuracy and robustness of the test set, incorporating expert-annotated data into the training set could bolster its reliability. Third, while we framed the classification of nuanced levels of participation in collective action as a multi-class task, it could alternatively have been approached as a multi-label task, given that multiple levels may apply to a single text. Last, our experiments are almost entirely based on Reddit data, which calls for more extensive investigation of the generality of our approach across platforms and contexts.
Despite its limitations, the classification method we propose holds promise for various applications that merit future research. First, it can enhance community targeting for activism campaigns by identifying broader audiences receptive to collective action causes on large-scale social media platforms, beyond traditional niche “activist” communities. This is particularly relevant on Reddit, where political discussions often occur in non-political subreddits (Rajadesingan, Budak, and Resnick 2021). Second, our classifiers can generate a large-scale and granular ground truth, which is invaluable for training and testing opinion dynamics models (Lenti, Monti, and De Francisci Morales 2024). These models, critical for understanding collective action dynamics, are often limited by insufficient ground truth data on opinion shifts and the social interactions driving them. Our classifiers address this by identifying relevant interactions indicative of a certain inclination towards collective action and their timing. Finally, applying our tool to online conversations relevant to social movements and activism organizations could enhance our understanding of the interplay between online activity and offline engagement among crowds (Smith et al. 2023).
Related Work
The study of social movements and collective action has long been of interest to social scientists, particularly in identifying participation predictors such as injustice, efficacy, and identity (Van Zomeren, Postmes, and Spears 2008; Furlong and Vignoles 2021; Thomas et al. 2022). In online contexts, these theories have been operationalized using dictionary-based approaches (Gulliver, Fielding, and Louis 2021; Brown, Lowery, and Smith 2022) and network-based enhancements (Erseghe et al. 2023). In a similar context, Smith, McGarty, and Thomas (2018) developed a dictionary of collective action terms which can be used to compute the percentage of words in text samples that align with these terms (Pera and Aiello 2024). However, these methods struggle to capture nuanced expressions of collective action and are limited by their reliance on fixed vocabularies, partial handling of word variations, and biases introduced by document length.
In contrast, classifiers based on text embeddings have been used to extract social dimensions such as emotions, empathy (Choi et al. 2020, 2023), and specific stages of collective action like calls-to-action (Rogers, Kovaleva, and Rumshisky 2019). While these methods provide greater flexibility, their performance is often constrained by their generalizability across contexts.
The advent of Large Language Models (LLMs) has created new possibilities for computational social science, especially in data-scarce scenarios (Møller and Aiello 2024). While LLMs offer broad utility, their effectiveness relative to fine-tuned classifiers depends on the task and domain, with trade-offs in annotation and computational costs. In some cases, simpler fine-tuned classifiers outperform LLMs in specific tasks (Ziems et al. 2024), underscoring the importance of task-specific model selection.
Reddit has been a valuable source for analyzing group-specific characteristics, such as mental health (Kim et al. 2023) and political leanings (Hofmann, Schütze, and Pierrehumbert 2022). However, individual engagement in activism through textual traces remains underexplored. Studies often use proxies like subreddit involvement to measure activism (Lenti et al. 2024), overlooking comment-level commitments to collective action.
Appendix A Data and Model Training
A.1 Subreddits on Activism
Table A1: List of subreddits included in the set of reference for training and testing.
Subreddit N. subscribers Creation date
antiwork 2,814,235 2013-08-14
EndFPTP 11,481 2016-07-29
neurodiversity 75,768 2012-04-27
Dyslexia 27,494 2010-11-13
AmIFreeToGo 52,296 2012-10-25
FreeSpeech 51,552 2009-02-15
prochoice 44,147 2010-01-15
jillstein 6,654 2011-04-20
LockdownSkepticism 54,693 2020-03-25
publicdefenders 10,606 2012-10-30
SyrianRebels 5,224 2016-07-30
TwoXChromosomes 13,666,599 2009-07-16
againstmensrights 24,949 2011-07-20
europrivacy 20,212 2014-03-09
ADHDers 21,574 2020-06-01
Vegetarianism 45,310 2008-05-17
ShitLiberalsSay 184,152 2012-07-24
WhiteRights - 2010-03-01
runaway 13,439 2011-11-16
MayDayStrike 46,993 2022-01-05
Egalitarianism 11,267 2011-04-05
terfisaslur - 2017-02-02
humanrights 14,471 2009-04-16
antiwar 13,248 2008-08-01
troubledteens 43,742 2011-03-26
AnimalRights 25,706 2008-04-15
HerpesCureResearch 24,248 2020-04-10
FeMRADebates 17,302 2013-08-06
Intactivism 7,086 2009-03-07
evolutionReddit 15,403 2012-01-23
IronFrontUSA 37,519 2018-05-20
water 43,907 2008-07-10
Political_Revolution 171,161 2016-02-03
activism - 2008-03-09
VeganActivism 21,081 2013-01-01
climate 189,199 2008-05-07
ndp 32,909 2008-09-08
israelexposed 54,630 2009-09-30
MensRights 359,895 2008-03-19
postnationalist 3,644 2013-08-27
badgovnofreedom 4,277 2013-09-06
EvolveSustain - 2012-06-04
Table A1 contains the list of subreddits that are considered as reference for the definition of the training and test sets, together with the number of subscribers as of April 2024 (dataset creation date). Subreddits with a missing number of subscribers are either banned or private but their data before their change in status can still be retrieved.
When filtering the set of top 40,000 subreddits by activism-related keywords, we excluded subreddits mainly containing non-English content (e.g. r/écologie), generic communities related to countries or continents (e.g. r/unitedstatesofindia), communities discussing artists or public figures not directly related to activism (e.g. r/chancetherapper), subreddits discussing economic theories not specifically related to activism (e.g. r/mmt_economics), communities discussing games, movies or books out of the context of activism (e.g. r/501st), subreddits that went private as a protest to Reddit API changes in 2023 and for which it is not possible to retrieve the topical focus from the name (e.g. r/charcoal), debate-focused subreddits (e.g. r/debateavegan), subreddits in which “rights” are mentioned only in connection to copyright (e.g. r/freesounds), rental or unemployment rights (e.g. r/loveforlandlords, r/texasunemployment) but not specifically related to the context of activism, subreddits in which drugs and their use are discussed and are not strictly related to activism on the topics (e.g. r/stoners), communities discussing bots rights (e.g. r/botsrights).
A.2 Crowdsourcing Task
We report a screenshot of an example annotation on MTurk in Figure A1.
Figure A1: Example screenshot of the annotation task on MTurk.
Annotation Task Instructions
Given the social media comment, which is the most appropriate label to represent the expressed level of participation in collective action? Refer to the ’Instructions’ panel.
Instructions
We are interested in collective action expressed in social media comments. Collective action can refer to a broad range of topics, such as environment, civil rights and mental health, and can reflect a spectrum of stances. Consider the given labels, roughly defined on a scale of increasing level of participation in collective action:
  • Problem-solution: the comment identifies an issue relevant to collective action as problematic, potentially assigning blame to the responsible and proposing a view on how to address the problem.
  • Call-to-action: the comment attempts to convince others to join or support a cause (e.g. a “call to arms”).
  • Intention: the comment expresses the author’s willingness or interest to act.
  • Execution: the comment directly reports the author’s involvement in some form of action.
  • None: None of the above.
Annotate the social media comment below according to the labels. To ease the labeling process, if multiple labels apply please select the highest on the scale (e.g. if both “Intention” and “Execution” apply, select “Execution”). Please press ’More Instructions’ and read an example for each label before annotating. Any suspected fraud or abuse will result in forfeiture of compensation.
Examples for Each Label
  • Problem-solution: “The decrease gender violence, one should start from school education and then offer continuous support through the workplace.”
  • Call-to-action: “The wealth gap is widening every day, and it’s time for action. Join us in calling for economic justice by attending the rally downtown this Saturday.”
  • Intention: “I wish to do my part in fighting racial injustice and join some community meetings.”
  • Execution: “Just got back from volunteering at the local food bank. Small actions can make a difference.”
A.3 Annotation Codebook
We are annotating excerpts from social media comments and looking for instances of participation in collective action, expressed at different levels. In some cases, background knowledge (e.g. veganism being a social movement that advocates for the end of animal suffering and/or limiting resources for animal products to contrast climate change) might be useful for the annotations.
We define a collective action problem as:
  • Human-generated issue. An issue caused by social, political or economic factors faced by group of people (or an individual as an example of a symptom within a broader issue). The issue of electing a political candidate is considered a human-generated issue. Note that such definition can be slightly stretched to include animal rights. The problem isn’t a natural disaster or accident but is the result of human actions, decisions or inaction (e.g. flooding is not a collective action problem, lack of government support to flooding is). Examples:
    • Collective action problem: “All of the depression treatment was basically just ’it’s all in your head.’ They would just force you to do shit and call it treatment. If I did not participate in a module they convinced me that I was fucking up and prolonging my treatment.” The message reports an individual example of a pattern of a systemic problem, making it a collective action problem. Note that the possibility of extending the specific example to a broader problem should be present (in this case “they would just force you to do shit” helps generalizing the issue).
    • Not a collective action problem: “You are actively harmful to people because of your refusal to listen to them or believe that their suffering could be worse than yours. You say you don’t believe that delusions don’t exist in response to me talking about self-harm that happened due to delusion. You suggest that I go off the medication that makes it so I don’t kill myself to try to make my own.” This focuses on an interpersonal conflict with one person invalidating the other’s mental health experience, without the link to a systemic collective issue.
    • Not a collective action problem: “Don’t want that to happen to you? Don’t drink and drive. And if you do, then submit to the tests as you admitted to do when you obtained a drivers license.” This message does not identify any larger systemic issue that would require collective action. It refers to “that” without specifying what the issue is, and it does not explicitly specify how drinking and driving would impact a broader group.
  • Current. The problem is not historical or set in the past or, if it is, it should still be ongoing or there should be a clear identification of how it has impact on the present or on current or future actions that can be taken to mitigate it (e.g. Vietnam war faults is not a collective action problem, the discrimination of black people that started with slavery in the US is).
  • With opportunity for collective action. The issue can potentially be mitigated through group action (e.g. public advocacy, policy change, collective efforts, protests, social mobilization) or collective individual action (e.g. recycling).
  • Involves shared responsibility. Individuals in the in-group or even out-group members can take some form of action that contributes to alleviating the impacts of the issue.
The message needs to clearly refer to a collective action problem, i.e. coders should be able to identify and report the issue from the message itself, or determine the collective action problem as a broader expansion of a symptom described in the message. If the messages does not clearly report a collective action problem but refers to a social movement that is commonly known to be related to collective action problems (e.g. veganism) or describes the author as an activist or willing to become an activist, given the semantic charge of the concepts the message is still valid as defined within a collective action problem umbrella. In any other case, the message should be classified as None.
Coders should be particularly careful when the message contains parts of discourse simply reported from others: if the author adds information on top of the reported speech that addition might change the direction of the reported message.
Note that more than one commitment level might apply. When in doubt, always choose the higher ranked level (i.e. if “Intention” and “Execution” both apply, choose “Execution”).
Problem-solution
The comment highlights the collective action problem described within the message as actually problematic, potentially blaming those responsibilities and/or suggesting a way to fix it. Note that the suggestion of solutions could be expressed both in terms of collective or individual actions or steps to be taken, and will be generally about outlining steps to be taken rather than motivating others to join a cause. Opinions themselves can help identifying an issue as problematic, but stances or opinions without a clear identification of a collective action problem as actually problematic do not count. A similar argument can be considered for simple reports of facts that mention a collective action problem without identifying the issue as problematic. e.g. “I support reproduction rights, and I am aware that everyone has their opinion on it.” discusses a collective action problem, that is reproduction rights, but it does not actually identify it as a problem in the message and thus it does not count as a “Problem-solution” message.
A message referring to a collective action problem can be seen as identifying the issue as problematic even through its tone, for instance if emotionally charged and critical. e.g. “I think there’s a couple of problems with what you are saying. One gender is shipped off to fight and die. The other stays home.” uses a quite strong language to talk about the collective action problem of gender roles. e.g. “Men work more than women, it’s a fact.” relates to the gender roles collective action problem but does not identify it as a problem.
There is a subtle difference between proposing solution and calling for action (see next class), but the main difference lies in the sense of urgency, need for immediate action and giving motivation for others to join a cause that the Call-to-action class expresses. Usually, message that use the imperative tense should be considered as leaning towards Call-to-action, and messages using “X should do Y” more towards Problem-solution, but the contextual difference can identify the message in one category or the other. e.g. “If you want to help with the discrimination issue, you should contact one of the organizations listed in the submission post.” is considered a proposed solution and thus part of the “Problem-solution” class. e.g. “Join one of the organizations listed in the submission post to help with the discrimination issue.” implies a sense of higher urgency and would be part of the “Call-to-action” class. While the semantic meaning of the message is the same as the one above, the framing makes the difference.
Call-to-action
The comment tries to convince the audience to take a specific action (advocating, organizing, protesting, boycotting, legal action, but also “standing strong” or “support”) in response to a perceived issue within a collective action problem or become activists and to engage others in the action itself by placing an emphasis on collective Participation in real-time efforts. The proposed action could be immediate or long-term and is framed as necessary for addressing a collective problem, issue, or injustice. This action is often presented as a moral or ethical imperative, described with a sense of urgency, and often calls for a collective involvement.
If the message contains a Call-to-action but there is no collective action problem or mention of activism/movement that could be solved or tackled with the action, for instance because there is simply an expression of stance or opinion by the commenter, then the message should not be coded as part of the Call-to-action class. e.g. “I think the Republicans are ruining the US. We need to stop them now.” expresses a sense of urgency and contains a Call-to-action but it does not refer to any collective action problem. Since we are not interested in extracting frames in general but specific participation in collective action that consider our definition of collective action problem, the message would not be coded as containing the Call-to-action class.
Intention
The commenter shares their own desire to do something or give help to mitigate a collective action problem or be involved as an activist. Note that the collective action problem and/or the reference of an activist initiative (or a political/social movement) should be clear. Otherwise, the message can be considered as Intentional in its framing but it is hard to distinguish it from mere Intention to act in a general context. e.g. “I’ll participate if I can choose my conditions.” is a message that is too generic and cannot be associated to any collective action problem to be solved or activist movement.
Not that conditional Intention still counts as part of the Intention class. e.g. “I’ll donate to the men’s rights movement if it is going to a cause that is worth it.”
Execution
The commenter is describing their personal experience taking tangible actions to mitigate a collective action problem or being an activist. Both online and offline actions apply. Note that simply stating that one has been “fighting”, “standing for”, “supporting” does not qualify as Execution because it could be simply interpreted as interest on the issue rather than tangible actions taken to solve it.
Execution expressed without a clear definition of a collective action problem, a movement or being an activist is not enough for a message to be classified in the Execution class. e.g. “I donated to them and would do it again.” it is not clear in which context we are talking about donations, and it might as well be that it is not related to a collective action problem at all. If a political or social movement hint would be included, then the message could be coded as containing the Execution class. Note that expressions such as “I am a climate change activist” count as part of the Execution class even if no action is clearly stated. Being an activist is charged enough for the expression to be considered Execution.
None
The comment does not fit into one of the previous categories; its purpose isn’t clear nor relevant to a collective action problem.
Table B2: Performance metrics for the multi-class classification task performed directly.
Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
Centr. 0.37 0.13 0.04 0.05 0.38 0.2 0.36
BERT CS 0.36 0.51 0.07 0.09 0.41 0.29 0.40
Syn 0.4 0.49 0.17 0.28 0.62 0.39 0.56
ZS No def 0.35 0.51 0.04 0.36 0.38 0.33 0.37
Def 0.38 0.41 0.12 0.27 0.44 0.32 0.42
SFT (Def) CS 0.4 0.61 0.08 0.33 0.5 0.39 0.48
Ext. 0.34 0.41 0 0.1 0.49 0.27 0.45
Ext. + Syn I/E 0.34 0.34 0.11 0.33 0.51 0.33 0.46
CS + Syn I/E 0.39 0.66 0.2 0.47 0.55 0.45 0.52
CS + Syn A 0.39 0.56 0.17 0.48 0.56 0.43 0.52
DPO (Def) CS 0.41 0.48 0 0.4 0.63 0.39 0.57
Ext. 0.33 0.22 0.21 0 0.48 0.25 0.43
Ext. + Syn I/E 0.35 0.28 0.29 0.32 0.66 0.38 0.58
CS + Syn I/E 0.37 0.54 0.2 0.37 0.4 0.38 0.4
CS + Syn A 0.39 0 0.23 0.42 0.51 0.31 0.46
A.4 Modeling Details
For all fine-tuned LLM models, we set aside 10% of the training data as a validation set.
Both RoBERTa classification models, for the binary and the multi-class tasks, are trained for 30 epochs, with a learning rate of $4e^{-5}$ and a train batch size of 16. Given the presence of class imbalance, the model is adjusted with weights computed as $\frac{tot\_sample}{2*[samples\_class_{0},...,samples\_class_{n}]}$. All other parameters are left at their default values and details can be checked in the shared code.
All Llama3 fine-tuned models are used with a temperature set to 0.01. The SFT approach is trained for 20 epochs and the DPO approach for 10 epochs. Both models are set with a learning rate of $2e^{-4}$, a weight decay of 0.001, a max_grad_norm of 0.3, and a warmup ratio of 0.03. We set the QLoRa $\alpha$ to 16, the dropout rate to 0.05, and the matrix rank to 8. The DPO is trained with 2 randomly chosen examples of wrong responses for the multi-class task.
For the zero-shot LLM, we used enhanced AI prompting. The prompt of reference is reported in Appendix C. The starting point labels definitions are those used in the crowdsourcing annotation campaign and the following is the list of generated labels definitions:
  • Problem-Solution: The comment highlights an issue and possibly suggests a way to fix it, often naming those responsible.
  • Call-to-Action: The comment asks readers to take part in a specific activity, effort, or movement.
  • Intention: The commenter shares their own desire to do something or be involved in solving a particular issue.
  • Execution: The commenter is describing their personal experience taking direct actions towards a common goal.
  • None: The comment doesn’t fit into one of these categories; its purpose isn’t clear or relevant to collective action.
A.5 Licenses
The Llama-3-8B Instruct model employed for the zero-shot, SFT and DPO approaches is an open-source LLM released under a commercial use license [^2] https://www.llama.com/llama3/license/. The S-BERT model for the extraction of sentence embeddings is licensed under the Apache License 2.0., while the RoBERTa base model used for building the BERT-based classifier and the BERTopic model are licensed under a MIT license. The collective action dictionary by Smith, McGarty, and Thomas (2018) is publicly available and so are the stance detection dataset on climate change created by (Luo, Card, and Jurafsky 2020) and the repository of UK parliamentary debates transcripts (MySociety 2024). The community embeddings and social dimensions similarity data produced by Waller and Anderson (2021) are licensed under the CC-BY 4.0 license.
Appendix B Results
B.1 Additional Experiments on Test Set
Table B2 reports the performance of the classification approaches applied directly to the multi-class task for identifying the nuances of participation in collective action.
Table B3 reports the performance of the LLM fine-tuning-based classification approaches that do not consider a definition of what a collective action problem is in the prompt. Note that both results of the direct application to the multi-class task and results of the multi-class task as a second step of the layered framework are reported.
Table B3: Performance metrics for the multi-class classification task for the LLM fine-tuning approaches. This table considers the multi-class task performed both directly and as a second step of the layered approach, without considering the definition of what a collective action problem is in the prompt.
Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
SFT (direct) CS 0.42 0.49 0.18 0.31 0.6 0.4 0.55
Ext. 0.35 0.34 0.1 0.1 0.5 0.28 0.45
Ext. + Syn I/E 0.35 0.26 0.14 0.14 0.53 0.29 0.48
CS + Syn I/E 0.4 0.6 0.2 0.43 0.62 0.45 0.57
CS + Syn A 0.4 0.57 0.15 0.45 0.57 0.43 0.53
DPO (direct) CS 0.38 0.38 0.07 0.3 0.56 0.34 0.51
Ext. 0.34 0.32 0.24 0.19 0.49 0.32 0.45
Ext. + Syn I/E 0.34 0.3 0.22 0.22 0.58 0.33 0.51
CS + Syn I/E 0.4 0.39 0.17 0.39 0.59 0.39 0.54
CS + Syn A 0.38 0.35 0.1 0.47 0.52 0.36 0.48
SFT (layer) CS 0.36 0.45 0 0.24 0.78 0.37 0.67
Ext. 0.36 0.48 0.11 0.12 0.78 0.37 0.67
Ext. + Syn I/E 0.37 0.45 0.11 0.35 0.78 0.41 0.68
CS + Syn I/E 0.36 0.56 0.19 0.38 0.78 0.46 0.68
CS + Syn A 0.36 0.6 0.38 0.48 0.78 0.52 0.69
DPO (layer) CS 0.38 0.48 0.1 0.31 0.78 0.41 0.41
Ext. 0.34 0.26 0.12 0 0.78 0.30 0.30
Ext. + Syn I/E 0.36 0.33 0.1 0.33 0.78 0.38 0.38
CS + Syn I/E 0.35 0.53 0.13 0.45 0.78 0.45 0.45
CS + Syn A 0.35 0.64 0.24 0.42 0.78 0.49 0.48
B.2 Additional Details on Topic Modeling
Table B4 shows the complete configuration of topics identified in the set involving one thread from r/intj, one thread from r/wallstreetbets, and one thread from r/IronFrontUSA. Since topics 4-8 are quite small in size and can be either considered noise or be attributed to r/intj (topic 7, topic embedding similarity with topic 0 about covid and vaccines: 0.527) and r/IronFrontUSA (topic 8, topic embedding similarity with topic 1 about voting: 0.628), we focused on topics 0-3 for the sake of simplicity in the visualization reported in the main text. For completeness, we report the same visualization considering all topics in Figure B2.
Table B4: List of topics identified by applying BERTopic on the reference dataset for topic modeling.
Topic n. Topic keywords N. comments
-1 internet, say, thats, yes 213
0 covid, effects, vaccine, risk 332
1 voting, vote, democrats, power 165
2 science, im, read, intj 141
3 robinhood, money, gme, app 105
4 dick, twist, it, fuck 31
5 this, fr, rarity, baaaa 25
6 yes, shit, sure, no 21
7 booster, shot, soon, yes 16
8 government, white, targeted, person 10
Figure B2: Complete visual representation of the topic modeling application, resulting in 9 topics (plus noise).
Appendix C Prompts
Note that, for the Labels Descriptions Generations prompt, the label definitions to be re-written by the LLM are those presented to crowdworkers (see Section A.2).
Control Samples Generation/Data Augmentation Labels Definitions Generations LLM Binary Task (no def) LLM Binary Task (def) LLM Multi-class Task (no def) LLM Multi-class Task (def)

### v2

@@
 While some online movements fail to yield tangible outcomes, in multiple occasions the Social Web has demonstrated its transformative potential in mobilizing collective action (Gerbaudo 2012) to achieve various political, societal, and economic goals (Yasseri et al.
-2016; Chiovaro et al.
-2021; Boulianne, Lalancette, and Ilkiw 2020; Lucchini et al.
+2016; Boulianne and Theocharis 2020; Lucchini et al.
 2022).
@@
 The operationalization of social science theories relevant to collective participation, often overlooked in current literature on online behavior change (Pera, Morales, and Aiello 2023), could drive progress in this respect and support real-world applications.
-To address these challenges, we combined insights from literature on mobilization and activism with modern Natural Language Processing (NLP) techniques to develop open-source text classifiers to detect expressions of participation in collective action from social media posts [^1] Data and code are available at https://tinyurl.com/4tj57wfk.
+To address these challenges, we combined insights from literature on mobilization and activism with modern Natural Language Processing (NLP) techniques to develop open-source text classifiers to detect expressions of participation in collective action from social media posts [^1] Data and code are available at https://github.com/ariannap13/extract˙collective˙action.
 Models are shared at https://huggingface.co/ariannap22..
@@
 We found that small models can reliably quantify participation in collective action, often rivaling Large Language Models (LLMs).
-When applied to the binary classification task of identifying any expression of participation, the best BERT-based model achieved a macro F1 score of $0.71$, close to the $0.77$ of the best-performing LLM approach.
+When applied to the binary classification task of identifying any expression of participation, the best BERT-based model achieved a macro F1 score of $0.65$, close to the $0.71$ of the best-performing LLM approach.
 In the multi-class classification of participation levels, the best BERT model obtained a macro F1 score of $0.44$, not far from the $0.52$ obtained by the best fine-tuned LLM.
@@
 Such insights demonstrate the potential of our approach to considerably broaden the scope of studies of collective action in online settings.
+Related Work
+Social Movements and Collective Action
+The study of social movements and collective action has long been of interest to social scientists, particularly in identifying participation predictors such as injustice, efficacy, and identity (Van Zomeren, Postmes, and Spears 2008; Furlong and Vignoles 2021; Thomas et al.
+2022).
+In online contexts, these theories have been operationalized using dictionary-based approaches (Gulliver, Fielding, and Louis 2021; Brown, Lowery, and Smith 2022) and network-based enhancements (Erseghe et al.
+2023).
+In a similar context, Smith, McGarty, and Thomas (2018) developed a dictionary of collective action terms which can be used to compute the percentage of words in text samples that align with these terms (Pera and Aiello 2024).
+However, these methods struggle to capture nuanced expressions of collective action and are limited by their reliance on fixed vocabularies, partial handling of word variations, and biases introduced by document length.
+Language Models in Text Classification
+Classifiers based on text embeddings have been used to extract social dimensions such as emotions, empathy (Choi et al.
+2020, 2023), and specific stages of collective action like calls-to-action (Rogers, Kovaleva, and Rumshisky 2019).
+While these methods provide greater flexibility, their performance is often constrained by their generalizability across contexts and scopes.
+The advent of Large Language Models (LLMs) has created new possibilities for computational social science, especially in data-scarce scenarios (Møller and Aiello 2024).
+Prior work has explored their utility in tasks such as stance detection (Lan et al.
+2024), hate speech identification (Kumarage, Bhattacharjee, and Garland 2024), and misinformation detection (Huang et al.
+2025), often stressing the importance of hybrid approaches in such tasks.
+However, less attention has been paid to the ability of LLMs to classify user participation in collective action, a task requiring a nuanced understanding of social movement dynamics.
+Unlike previous studies that focus on a single classification paradigm, our work systematically compares dictionary-based, centroid-based, BERT-based, and LLM-based classifiers to assess their effectiveness in capturing participation signals.
+This comparative approach allows us to evaluate generalizability across different modeling strategies, shedding light on the trade-offs between interpretability, accuracy, and computational cost in collective action classification.
+Specifically, while LLMs offer broad utility, their effectiveness relative to fine-tuned classifiers depends on the task and domain, with trade-offs in annotation and computational costs.
+In some cases, simpler fine-tuned classifiers outperform LLMs in specific tasks (Ziems et al.
+2024), underscoring the importance of task-specific model selection.
+Studying Activism on Reddit
+Reddit has been a valuable source for analyzing group-specific characteristics, such as mental health (Kim et al.
+2023) and political leanings (Hofmann, Schütze, and Pierrehumbert 2022).
+However, individual engagement in activism through textual traces remains underexplored.
+Existing studies often use proxies like subreddit involvement to measure activism (Lenti et al.
+2025), overlooking comment-level commitments to collective action.
+In this work, we address this limitation by focusing on the extraction of comment-level participation in collective action, offering a new perspective on how online communities express engagement in activism.
 Theoretical Framework
@@
 We began our analysis by considering the top 40,000 subreddits ranked by all-time subscriber count, from which we identified a subset of 73 subreddits that contained the terms “activism,” “activist,” or “rights” in their titles or descriptions.
+The identification of these subreddits was performed automatically by searching for the presence of these seed terms, which were selected for their direct relevance to the topic of collective action.
 Following a manual review, we selected 42 subreddits pertinent to our study’s scope, containing a total of 120 million comments posted from the creation of these subreddits through December 2023 (see Section A.1 in Appendix A for the exclusion criteria and the list of selected subreddits).
@@
 The resulting annotated dataset consists of 369 samples: 100 labeled as None and 269 as expressing Participation in collective action, with 202 labeled as Problem-solution, 44 as Call-to-action, 14 as Intention, and 9 as Execution.
+Later on, we will refer to this dataset as CS.
 Data Augmentation
-To address data sparsity in the annotated data, particularly in classes representing higher levels of participation, we applied two data augmentation techniques (Table 1).
+To address data sparsity in the annotated data, particularly in classes representing higher levels of participation, we applied two data augmentation techniques.
 The first approach involved generating synthetic samples using Llama3 8B Instruct (Møller et al.
@@
 We also created a variation focusing solely on the minority classes Intention and Execution (Syn I/E).
-All generated samples were manually reviewed for accuracy.
 The generation prompt is provided in Appendix C.
+All generated samples were manually reviewed for accuracy using a pipeline based on the validity criteria from the prompt for each anchor example (e.g., semantic similarity, structure, meaning, intent, and key details).
+A table was created to binary-check the samples against these criteria.
 The second method used Reddit’s structural features.
@@
 The concatenation of the crowdsourced annotated dataset and this Reddit-informed extension is referred to as Ext.
-Table 1: N. of participation labels by augmentation type.
+Table 1 summarizes the label volumes for the crowdsourced dataset (CS), the synthetic augmentation for minority classes (CS + Syn I/E), the synthetic augmentation for all classes (CS + Syn A), the Reddit-informed extension (Ext), and the combination of both Reddit extension and synthetic augmentation for minority classes (Ext + Syn I/E).
 Type P-Sol C-Act Intent Exec None
@@
 Ext + Syn I/E 553 175 285 189 285
+Table 1: N. of participation labels by augmentation type.
 Test Set
@@
 For binary classification, we used the dictionary of collective action words (Smith, McGarty, and Thomas 2018) to calculate the fraction of such words in the text.
-We then selected the threshold maximizing the difference between the True Positive Rate (TPR) and False Positive Rate (FPR) on the training set.
+We then selected the threshold maximizing the difference between the True Positive Rate (TPR) and False Positive Rate (FPR) on the training set, prioritizing a class-balanced decision boundary to avoid under-detecting instances of “Participation”.
+While we initially considered maximizing the F1-score, we found that it tended to favor the majority class, leading to a threshold that disproportionately benefited the more frequent “None” class.
+In contrast, maximizing $TPR-FPR$ allowed for a more balanced threshold, ensuring better detection of “Participation” cases without bias towards the majority class.
 Centroid classifier (Centr)
@@
 Table 2 reports the performance of the binary classification task–the first step in our layered approach–measured in terms of F1 scores.
+Details on precision and recall values can be found in Table  B2 in Appendix B.
 The best-performing models were the zero-shot Llama3 model (both with and without the collective action problem definition in the prompt) and the RoBERTa classifier trained on the synthetically augmented training set.
@@
 2023), the higher computational costs during inference and the complexity of hosting the models locally, we decided to select the RoBERTa model as the starting point for the multi-class step of the layered approach.
-Table 2: F1 scores for the binary classification task. The three best-performing models are in bold.
-Method Train. set Part. None Macro Avg. Weight. Avg.
+Method Train. set Part. None Macro Avg. Weigh. Avg.
 Dict. 0.39 0.51 0.45 0.48
@@
 BERT CS 0.44 0.28 0.36 0.32
-Syn A 0.51 0.78 0.65 0.71
+CS + Syn A 0.51 0.78 0.65 0.71
 ZS no def 0.56 0.84 0.70 0.76
@@
 CS (def) 0.48 0.47 0.47 0.47
-Syn A (no def) 0.50 0.59 0.54 0.57
-Syn A (def) 0.51 0.59 0.55 0.57
-DPO Syn A (def) 0.47 0.42 0.44 0.43
-Table 3: Performance metrics for the multi-class classification task as a second step of the layered approach. The three best-performing models are in bold.
+CS + Syn A (no def) 0.50 0.59 0.54 0.57
+CS + Syn A (def) 0.51 0.59 0.55 0.57
+DPO CS + Syn A (def) 0.47 0.42 0.44 0.43
+Table 2: F1 scores for the binary classification task, with the top three models in bold.
+CS denotes the crowdsourced dataset, CS + Syn A indicates synthetic augmentation for all classes, and def and no def refer to whether the characteristics of a collective action problem are defined in the LLM prompt or not.
 Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
@@
 BERT CS 0.36 0.56 0.11 0.25 0.78 0.41 0.68
-Syn A 0.36 0.62 0.17 0.29 0.78 0.44 0.68
-ZS no def 0.34 0.54 0.09 0.46 0.78 0.44 0.67
+CS + Syn A 0.36 0.62 0.17 0.29 0.78 0.44 0.68
+ZS no def 0.34 0.54 0.09 0.46 0.79 0.44 0.68
 def 0.38 0.39 0.09 0.36 0.78 0.40 0.67
@@
 CS + Syn A 0.38 0.58 0.29 0.5 0.78 0.51 0.69
+Table 3: F1 scores for the multi-class classification task as a second step of the layered approach, with the top three models in bold.
+CS denotes the crowdsourced dataset, CS + Syn I/E indicates synthetic augmentation for minority classes, CS + Syn A for all classes, Ext denotes the Reddit-informed extension, and Ext + Syn I/E defines the combination of both Reddit extension and synthetic augmentation for minority classes.
+All LLM prompts define a collective action problem.
 Multi-class Task
@@
 During testing, we focused on comments selected in the first classification stage by the BERT-based classifier trained on synthetically augmented data.
-Table 3 reports performances for the multi-class task, considering the inclusion of the collective action problem definition in the LLM prompts.
+Table 3 reports performances for the multi-class task in terms of F1 scores, considering the inclusion of the collective action problem definition in the LLM prompts.
+Details on precision and recall values can be found in Table  B3 in Appendix B.
 The SFT Llama3 model, trained on crowdworker-annotated data with synthetic augmentation for all classes except the majority class (CS+Syn A), achieved the best overall performance.
@@
 However, the SFT and DPO models consistently outperformed others, particularly in classifying the minority classes Intention and Execution.
-The ranking of models remained largely consistent whether or not the collective action problem definition was included in the LLM prompts or when models were applied directly to the multi-class task without a layered approach (see Appendix B, Tables B2 and B3).
+The ranking of models remained largely consistent whether or not the collective action problem definition was included in the LLM prompts or when models were applied directly to the multi-class task without a layered approach (see Appendix B, Tables B4 and B5).
 Validation
@@
 Figure 1: Validation of the proposed approach.
-(a-d) comparison with topic modeling; (e) comparison with stance detection; (f) impact on classification performance given by the use of collective action keywords to define training and test sets.
+(a-d) comparison with topic modeling; (e) comparison with stance detection – A for agree, D for disagree, N for neutral; (f) impact of collective action keywords on classification performance.
 Topic Modeling
@@
 To ensure our classifier does not rely solely on the presence of these words, we tested the classifier on four variations of the test set: (i) removal of action dictionary words, (ii) substitution of action dictionary words with random tokens from the model vocabulary, (iii) removal of random words, and (iv) substitution of random words with random tokens.
+This approach aligns with established practices in NLP for evaluating model robustness using adversarial datasets (Ebrahimi et al.
+2017; Ghaddar et al.
+2021).
 If collective action words influenced the classifier, we would expect different results for configurations (i) and (ii) compared to the original test set.
@@
 Our approach effectively distinguishes messages expressing participation from those that do not, either by being unrelated to collective action or not identifying the issues as problematic (Table 4).
-Table 4: Selected examples of parliamentary debates transcripts from the top and bottom 10 probabilities of class Participation in collective action.
+While this analysis is preliminary, it demonstrates the potential of our approach for structured political discourse, highlighting a promising avenue for future research on collective action detection in institutional settings.
 Text
@@
 3. How many prisoners have been released early under the end of custody supervised licence scheme since October 2023.
-Figure 2: Comparison of climate change-focused comment percentages and participation in collective action levels across subreddits, with x-axis values normalized using min-max scaling.
+Table 4: Selected examples of parliamentary debates transcripts from the top and bottom 10 probabilities of class Participation in collective action.
+Figure 2: Comparison of climate change-focused comment percentages and participation in collective action levels across subreddits, x-axis values normalized.
 Climate Action
@@
 2022; Lenti et al.
-2024).
+2025).
 In contrast, we show that subreddits selected based on keyword counts may not always align with actual participation in collective action.
@@
 2018; Lenti et al.
-2024).
+2025).
 However, such studies often rely on coarse proxies, such as subreddit participation as a stand-in for activism on Reddit (Lenti et al.
-2024).
+2025).
 Our classification pipeline presents a new opportunity to examine whether the socio-political dimensions estimated by state-of-the-art embedding methods correlate with direct expressions of participation in collective action, and whether these correlations align with theoretical expectations.
@@
 Notably, our BERT-based classifier offers an optimal balance between performance, computational efficiency, and cost-effectiveness across both layers of classification (RQ1).
+While the specific reasons behind the comparable performance of small language models and fine-tuned LLMs on our binary classification task remain unclear, similar findings have been reported in both NLP and Computational Social Science literature (Bucher and Martini 2024; Ziems et al.
+2024).
+This suggests that smaller, domain-specific models can rival larger models for certain tasks, and we consider this an important avenue for future research.
 In applying our tool to data derived from Reddit discussions, we observed that our classification yields results that are orthogonal to those produced by traditional topic modeling and stance detection algorithms.
@@
 Furthermore, our analysis indicates that discussions rich in expressions of participation in collective action are prevalent in forums characterized by more female, adult and left-leaning participants, in line with theoretical expectations (RQ3).
+Limitations and Future Work
 Our proposed method has some limitations.
-First, the volume of the annotated training set is relatively small.
-Although we have implemented data augmentation techniques to address this limitation, the collection of additional annotations could further enhance classification quality.
-Second, while the training set consists of crowdsourced annotations, our test set is annotated by domain experts.
+First, our focus on a limited number of seed terms—“activism,” “activist,” and “rights”—may have biased the selection of subreddits, potentially overlooking relevant but ideologically diverse communities.
+This could skew the dataset toward left-leaning discourse, affecting downstream applications.
+Future work could refine subreddit selection by incorporating additional terms or applying more systematic, politically balanced criteria.
+Second, the volume of the annotated training set is relatively small.
+While we have validated the augmented data and used it strategically to mitigate the cost of crowdsourcing annotations, its influence on model performance and potential biases warrants further investigation.
+This is particularly important given class imbalance and the presence of rare classes, underscoring the need for future research on their treatment and the development of tailored augmentation strategies to ensure robustness and reliability.
+Third, while the training set consists of crowdsourced annotations, our test set is annotated by domain experts.
 Although we prioritized the accuracy and robustness of the test set, incorporating expert-annotated data into the training set could bolster its reliability.
-Third, while we framed the classification of nuanced levels of participation in collective action as a multi-class task, it could alternatively have been approached as a multi-label task, given that multiple levels may apply to a single text.
-Last, our experiments are almost entirely based on Reddit data, which calls for more extensive investigation of the generality of our approach across platforms and contexts.
+Fourth, while we framed the classification of nuanced levels of participation in collective action as a multi-class task, it could alternatively have been approached as a multi-label task, given that multiple levels may apply to a single text.
+Fifth, our experiments are primarily based on Reddit data, with the training set exclusively sourced from Reddit.
+While this provides a sufficient foundation for an initial study, further research is needed to explore the generalizability of our approach across different platforms and contexts.
+Finally, while we provide an extensive analysis of model selection, configuration design, and training set variations, our exploration is not exhaustive.
+Future research could expand on these aspects by experimenting with additional modeling choices and training strategies.
+Potential Applications
 Despite its limitations, the classification method we propose holds promise for various applications that merit future research.
@@
 2023).
-Related Work
-The study of social movements and collective action has long been of interest to social scientists, particularly in identifying participation predictors such as injustice, efficacy, and identity (Van Zomeren, Postmes, and Spears 2008; Furlong and Vignoles 2021; Thomas et al.
-2022).
-In online contexts, these theories have been operationalized using dictionary-based approaches (Gulliver, Fielding, and Louis 2021; Brown, Lowery, and Smith 2022) and network-based enhancements (Erseghe et al.
-2023).
-In a similar context, Smith, McGarty, and Thomas (2018) developed a dictionary of collective action terms which can be used to compute the percentage of words in text samples that align with these terms (Pera and Aiello 2024).
-However, these methods struggle to capture nuanced expressions of collective action and are limited by their reliance on fixed vocabularies, partial handling of word variations, and biases introduced by document length.
-In contrast, classifiers based on text embeddings have been used to extract social dimensions such as emotions, empathy (Choi et al.
-2020, 2023), and specific stages of collective action like calls-to-action (Rogers, Kovaleva, and Rumshisky 2019).
-While these methods provide greater flexibility, their performance is often constrained by their generalizability across contexts.
-The advent of Large Language Models (LLMs) has created new possibilities for computational social science, especially in data-scarce scenarios (Møller and Aiello 2024).
-While LLMs offer broad utility, their effectiveness relative to fine-tuned classifiers depends on the task and domain, with trade-offs in annotation and computational costs.
-In some cases, simpler fine-tuned classifiers outperform LLMs in specific tasks (Ziems et al.
-2024), underscoring the importance of task-specific model selection.
-Reddit has been a valuable source for analyzing group-specific characteristics, such as mental health (Kim et al.
-2023) and political leanings (Hofmann, Schütze, and Pierrehumbert 2022).
-However, individual engagement in activism through textual traces remains underexplored.
-Studies often use proxies like subreddit involvement to measure activism (Lenti et al.
-2024), overlooking comment-level commitments to collective action.
+Acknowledgments
+We acknowledge the support from the Carlsberg Foundation through the COCOONS project (CF21-0432).
+The funder had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
 Appendix A Data and Model Training
 A.1 Subreddits on Activism
-Table A1: List of subreddits included in the set of reference for training and testing.
 Subreddit N. subscribers Creation date
@@
 EvolveSustain - 2012-06-04
+Table A1: List of subreddits included in the set of reference for training and testing.
 Table A1 contains the list of subreddits that are considered as reference for the definition of the training and test sets, together with the number of subscribers as of April 2024 (dataset creation date).
@@
 Figure A1: Example screenshot of the annotation task on MTurk.
-Annotation Task Instructions
-Given the social media comment, which is the most appropriate label to represent the expressed level of participation in collective action? Refer to the ’Instructions’ panel.
-Instructions
-We are interested in collective action expressed in social media comments.
-Collective action can refer to a broad range of topics, such as environment, civil rights and mental health, and can reflect a spectrum of stances.
-Consider the given labels, roughly defined on a scale of increasing level of participation in collective action:
-  • Problem-solution: the comment identifies an issue relevant to collective action as problematic, potentially assigning blame to the responsible and proposing a view on how to address the problem.
-  • Call-to-action: the comment attempts to convince others to join or support a cause (e.g. a “call to arms”).
-  • Intention: the comment expresses the author’s willingness or interest to act.
-  • Execution: the comment directly reports the author’s involvement in some form of action.
-  • None: None of the above.
-Annotate the social media comment below according to the labels.
-To ease the labeling process, if multiple labels apply please select the highest on the scale (e.g.
-if both “Intention” and “Execution” apply, select “Execution”).
-Please press ’More Instructions’ and read an example for each label before annotating.
-Any suspected fraud or abuse will result in forfeiture of compensation.
-Examples for Each Label
-  • Problem-solution: “The decrease gender violence, one should start from school education and then offer continuous support through the workplace.”
-  • Call-to-action: “The wealth gap is widening every day, and it’s time for action. Join us in calling for economic justice by attending the rally downtown this Saturday.”
-  • Intention: “I wish to do my part in fighting racial injustice and join some community meetings.”
-  • Execution: “Just got back from volunteering at the local food bank. Small actions can make a difference.”
+For annotation task instructions, see crowdsource_instructions.pdf in the GitHub repo https://github.com/ariannap13/extract˙collective˙action.
 A.3 Annotation Codebook
-We are annotating excerpts from social media comments and looking for instances of participation in collective action, expressed at different levels.
-In some cases, background knowledge (e.g.
-veganism being a social movement that advocates for the end of animal suffering and/or limiting resources for animal products to contrast climate change) might be useful for the annotations.
-We define a collective action problem as:
-  • Human-generated issue.
-An issue caused by social, political or economic factors faced by group of people (or an individual as an example of a symptom within a broader issue).
-The issue of electing a political candidate is considered a human-generated issue.
-Note that such definition can be slightly stretched to include animal rights.
-The problem isn’t a natural disaster or accident but is the result of human actions, decisions or inaction (e.g.
-flooding is not a collective action problem, lack of government support to flooding is).
-Examples:
-    • Collective action problem: “All of the depression treatment was basically just ’it’s all in your head.’ They would just force you to do shit and call it treatment.
-If I did not participate in a module they convinced me that I was fucking up and prolonging my treatment.” The message reports an individual example of a pattern of a systemic problem, making it a collective action problem.
-Note that the possibility of extending the specific example to a broader problem should be present (in this case “they would just force you to do shit” helps generalizing the issue).
-    • Not a collective action problem: “You are actively harmful to people because of your refusal to listen to them or believe that their suffering could be worse than yours.
-You say you don’t believe that delusions don’t exist in response to me talking about self-harm that happened due to delusion.
-You suggest that I go off the medication that makes it so I don’t kill myself to try to make my own.” This focuses on an interpersonal conflict with one person invalidating the other’s mental health experience, without the link to a systemic collective issue.
-    • Not a collective action problem: “Don’t want that to happen to you?
-Don’t drink and drive.
-And if you do, then submit to the tests as you admitted to do when you obtained a drivers license.” This message does not identify any larger systemic issue that would require collective action.
-It refers to “that” without specifying what the issue is, and it does not explicitly specify how drinking and driving would impact a broader group.
-  • Current.
-The problem is not historical or set in the past or, if it is, it should still be ongoing or there should be a clear identification of how it has impact on the present or on current or future actions that can be taken to mitigate it (e.g.
-Vietnam war faults is not a collective action problem, the discrimination of black people that started with slavery in the US is).
-  • With opportunity for collective action.
-The issue can potentially be mitigated through group action (e.g.
-public advocacy, policy change, collective efforts, protests, social mobilization) or collective individual action (e.g.
-recycling).
-  • Involves shared responsibility. Individuals in the in-group or even out-group members can take some form of action that contributes to alleviating the impacts of the issue.
-The message needs to clearly refer to a collective action problem, i.e.
-coders should be able to identify and report the issue from the message itself, or determine the collective action problem as a broader expansion of a symptom described in the message.
-If the messages does not clearly report a collective action problem but refers to a social movement that is commonly known to be related to collective action problems (e.g.
-veganism) or describes the author as an activist or willing to become an activist, given the semantic charge of the concepts the message is still valid as defined within a collective action problem umbrella.
-In any other case, the message should be classified as None.
-Coders should be particularly careful when the message contains parts of discourse simply reported from others: if the author adds information on top of the reported speech that addition might change the direction of the reported message.
-Note that more than one commitment level might apply. When in doubt, always choose the higher ranked level (i.e. if “Intention” and “Execution” both apply, choose “Execution”).
-Problem-solution
-The comment highlights the collective action problem described within the message as actually problematic, potentially blaming those responsibilities and/or suggesting a way to fix it.
-Note that the suggestion of solutions could be expressed both in terms of collective or individual actions or steps to be taken, and will be generally about outlining steps to be taken rather than motivating others to join a cause.
-Opinions themselves can help identifying an issue as problematic, but stances or opinions without a clear identification of a collective action problem as actually problematic do not count.
-A similar argument can be considered for simple reports of facts that mention a collective action problem without identifying the issue as problematic.
-e.g.
-“I support reproduction rights, and I am aware that everyone has their opinion on it.” discusses a collective action problem, that is reproduction rights, but it does not actually identify it as a problem in the message and thus it does not count as a “Problem-solution” message.
-A message referring to a collective action problem can be seen as identifying the issue as problematic even through its tone, for instance if emotionally charged and critical.
-e.g.
-“I think there’s a couple of problems with what you are saying.
-One gender is shipped off to fight and die.
-The other stays home.” uses a quite strong language to talk about the collective action problem of gender roles.
-e.g.
-“Men work more than women, it’s a fact.” relates to the gender roles collective action problem but does not identify it as a problem.
-There is a subtle difference between proposing solution and calling for action (see next class), but the main difference lies in the sense of urgency, need for immediate action and giving motivation for others to join a cause that the Call-to-action class expresses.
-Usually, message that use the imperative tense should be considered as leaning towards Call-to-action, and messages using “X should do Y” more towards Problem-solution, but the contextual difference can identify the message in one category or the other.
-e.g.
-“If you want to help with the discrimination issue, you should contact one of the organizations listed in the submission post.” is considered a proposed solution and thus part of the “Problem-solution” class.
-e.g.
-“Join one of the organizations listed in the submission post to help with the discrimination issue.” implies a sense of higher urgency and would be part of the “Call-to-action” class.
-While the semantic meaning of the message is the same as the one above, the framing makes the difference.
-Call-to-action
-The comment tries to convince the audience to take a specific action (advocating, organizing, protesting, boycotting, legal action, but also “standing strong” or “support”) in response to a perceived issue within a collective action problem or become activists and to engage others in the action itself by placing an emphasis on collective Participation in real-time efforts.
-The proposed action could be immediate or long-term and is framed as necessary for addressing a collective problem, issue, or injustice.
-This action is often presented as a moral or ethical imperative, described with a sense of urgency, and often calls for a collective involvement.
-If the message contains a Call-to-action but there is no collective action problem or mention of activism/movement that could be solved or tackled with the action, for instance because there is simply an expression of stance or opinion by the commenter, then the message should not be coded as part of the Call-to-action class.
-e.g.
-“I think the Republicans are ruining the US.
-We need to stop them now.” expresses a sense of urgency and contains a Call-to-action but it does not refer to any collective action problem.
-Since we are not interested in extracting frames in general but specific participation in collective action that consider our definition of collective action problem, the message would not be coded as containing the Call-to-action class.
-Intention
-The commenter shares their own desire to do something or give help to mitigate a collective action problem or be involved as an activist.
-Note that the collective action problem and/or the reference of an activist initiative (or a political/social movement) should be clear.
-Otherwise, the message can be considered as Intentional in its framing but it is hard to distinguish it from mere Intention to act in a general context.
-e.g.
-“I’ll participate if I can choose my conditions.” is a message that is too generic and cannot be associated to any collective action problem to be solved or activist movement.
-Not that conditional Intention still counts as part of the Intention class. e.g. “I’ll donate to the men’s rights movement if it is going to a cause that is worth it.”
-Execution
-The commenter is describing their personal experience taking tangible actions to mitigate a collective action problem or being an activist.
-Both online and offline actions apply.
-Note that simply stating that one has been “fighting”, “standing for”, “supporting” does not qualify as Execution because it could be simply interpreted as interest on the issue rather than tangible actions taken to solve it.
-Execution expressed without a clear definition of a collective action problem, a movement or being an activist is not enough for a message to be classified in the Execution class.
-e.g.
-“I donated to them and would do it again.” it is not clear in which context we are talking about donations, and it might as well be that it is not related to a collective action problem at all.
-If a political or social movement hint would be included, then the message could be coded as containing the Execution class.
-Note that expressions such as “I am a climate change activist” count as part of the Execution class even if no action is clearly stated.
-Being an activist is charged enough for the expression to be considered Execution.
-None
-The comment does not fit into one of the previous categories; its purpose isn’t clear nor relevant to a collective action problem.
-Table B2: Performance metrics for the multi-class classification task performed directly.
-Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
-Centr. 0.37 0.13 0.04 0.05 0.38 0.2 0.36
-BERT CS 0.36 0.51 0.07 0.09 0.41 0.29 0.40
-Syn 0.4 0.49 0.17 0.28 0.62 0.39 0.56
-ZS No def 0.35 0.51 0.04 0.36 0.38 0.33 0.37
-Def 0.38 0.41 0.12 0.27 0.44 0.32 0.42
-SFT (Def) CS 0.4 0.61 0.08 0.33 0.5 0.39 0.48
-Ext. 0.34 0.41 0 0.1 0.49 0.27 0.45
-Ext. + Syn I/E 0.34 0.34 0.11 0.33 0.51 0.33 0.46
-CS + Syn I/E 0.39 0.66 0.2 0.47 0.55 0.45 0.52
-CS + Syn A 0.39 0.56 0.17 0.48 0.56 0.43 0.52
-DPO (Def) CS 0.41 0.48 0 0.4 0.63 0.39 0.57
-Ext. 0.33 0.22 0.21 0 0.48 0.25 0.43
-Ext. + Syn I/E 0.35 0.28 0.29 0.32 0.66 0.38 0.58
-CS + Syn I/E 0.37 0.54 0.2 0.37 0.4 0.38 0.4
-CS + Syn A 0.39 0 0.23 0.42 0.51 0.31 0.46
+See annotation_codebook.pdf in the GitHub repo https://github.com/ariannap13/extract˙collective˙action.
+Method Train. set Participation None Macro Avg. Weight. Avg.
+P R P R P R P R
+Dict. 0.27 0.67 0.77 0.38 0.52 0.53 0.64 0.46
+Centr. 0.36 0.81 0.88 0.49 0.62 0.65 0.75 0.57
+BERT CS 0.29 0.96 0.93 0.17 0.61 0.56 0.76 0.37
+CS + Syn A 0.44 0.60 0.84 0.73 0.64 0.67 0.74 0.70
+ZS no def 0.53 0.58 0.85 0.82 0.69 0.70 0.77 0.76
+def 0.53 0.66 0.87 0.80 0.70 0.73 0.78 0.76
+SFT CS (no def) 0.32 0.88 0.89 0.35 0.61 0.61 0.74 0.49
+CS (def) 0.32 0.93 0.93 0.31 0.62 0.62 0.77 0.47
+CS + Syn A (no def) 0.35 0.86 0.90 0.44 0.62 0.65 0.76 0.55
+CS + Syn A (def) 0.36 0.90 0.93 0.43 0.64 0.67 0.78 0.55
+DPO CS + Syn A (def) 0.31 0.95 0.94 0.27 0.62 0.61 0.77 0.44
+Table B2: Precision (P) and recall (R) scores for the binary classification task.
 A.4 Modeling Details
@@
 Appendix B Results
-B.1 Additional Experiments on Test Set
+B.1 Additional Performance Metrics
+Table B2 reports the metrics of precision and recall for the binary task, while Table B3 them for the multi-class task.
+Method Train. Set Prob.-Sol. Call-to-A. Intention Execution None Macro Avg. Weigh. Avg.
+P R P R P R P R P R P R P R
+Centr. 0.33 0.45 0.19 0.13 0.03 0.09 0.08 0.15 0.84 0.73 0.29 0.31 0.69 0.63
+BERT CS 0.30 0.47 0.52 0.62 0.13 0.09 0.67 0.15 0.84 0.73 0.49 0.41 0.72 0.66
+CS + Syn A 0.30 0.43 0.63 0.62 0.12 0.27 0.27 0.31 0.84 0.73 0.43 0.47 0.71 0.66
+ZS no def 0.31 0.38 0.61 0.49 0.06 0.18 0.46 0.46 0.82 0.75 0.45 0.45 0.70 0.66
+def 0.32 0.49 0.83 0.26 0.08 0.09 0.44 0.31 0.85 0.78 0.50 0.38 0.73 0.68
+SFT (def) CS 0.30 0.48 0.56 0.56 0.20 0.09 0.50 0.38 0.84 0.73 0.48 0.45 0.72 0.67
+Ext 0.28 0.49 0.48 0.31 0.00 0.00 0.00 0.00 0.84 0.73 0.32 0.31 0.70 0.65
+Ext + Syn I/E 0.29 0.47 0.44 0.36 0.22 0.18 0.43 0.23 0.84 0.73 0.44 0.39 0.71 0.65
+CS + Syn I/E 0.30 0.45 0.54 0.56 0.15 0.18 0.45 0.38 0.84 0.73 0.46 0.46 0.71 0.66
+CS + Syn A 0.31 0.46 0.64 0.64 0.24 0.36 0.55 0.46 0.84 0.73 0.51 0.53 0.72 0.67
+DPO (def) CS 0.28 0.27 0.50 0.10 0.00 0.00 0.27 0.31 0.78 0.84 0.37 0.30 0.66 0.68
+Ext 0.28 0.46 0.47 0.41 0.00 0.00 0.00 0.00 0.84 0.74 0.32 0.32 0.70 0.65
+Ext + Syn I/E 0.27 0.45 0.35 0.21 0.00 0.00 0.33 0.08 0.82 0.74 0.36 0.29 0.68 0.64
+CS + Syn I/E 0.31 0.46 0.72 0.54 0.18 0.27 0.44 0.62 0.84 0.74 0.50 0.52 0.72 0.67
+CS + Syn A 0.33 0.45 0.50 0.69 0.30 0.27 0.42 0.62 0.84 0.74 0.48 0.55 0.72 0.68
+Table B3: Precision (P) and recall (R) metrics for the multi-class classification task as a second step of the layered approach.
+B.2 Additional Experiments on Test Set
 Table B4 reports the performance of the classification approaches applied directly to the multi-class task for identifying the nuances of participation in collective action.
+Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
+Centr. 0.37 0.13 0.04 0.05 0.38 0.2 0.36
+BERT CS 0.36 0.51 0.07 0.09 0.41 0.29 0.40
+CS + Syn A 0.4 0.49 0.17 0.28 0.62 0.39 0.56
+ZS No def 0.35 0.51 0.04 0.36 0.38 0.33 0.37
+Def 0.38 0.41 0.12 0.27 0.44 0.32 0.42
+SFT (def) CS 0.4 0.61 0.08 0.33 0.5 0.39 0.48
+Ext. 0.34 0.41 0 0.1 0.49 0.27 0.45
+Ext. + Syn I/E 0.34 0.34 0.11 0.33 0.51 0.33 0.46
+CS + Syn I/E 0.39 0.66 0.2 0.47 0.55 0.45 0.52
+CS + Syn A 0.39 0.56 0.17 0.48 0.56 0.43 0.52
+DPO (def) CS 0.41 0.48 0 0.4 0.63 0.39 0.57
+Ext. 0.33 0.22 0.21 0 0.48 0.25 0.43
+Ext. + Syn I/E 0.35 0.28 0.29 0.32 0.66 0.38 0.58
+CS + Syn I/E 0.37 0.54 0.2 0.37 0.4 0.38 0.4
+CS + Syn A 0.39 0 0.23 0.42 0.51 0.31 0.46
+Table B4: F1 scores for the multi-class classification task performed directly.
 Table B5 reports the performance of the LLM fine-tuning-based classification approaches that do not consider a definition of what a collective action problem is in the prompt.
 Note that both results of the direct application to the multi-class task and results of the multi-class task as a second step of the layered framework are reported.
-Table B3: Performance metrics for the multi-class classification task for the LLM fine-tuning approaches.
-This table considers the multi-class task performed both directly and as a second step of the layered approach, without considering the definition of what a collective action problem is in the prompt.
 Method Train. set Problem-solution Call-to-action Intention Execution None Macro Avg. Weight. Avg.
@@
 CS + Syn A 0.35 0.64 0.24 0.42 0.78 0.49 0.48
-B.2 Additional Details on Topic Modeling
+Table B5: F1 scores for the multi-class classification task for the LLM fine-tuning approaches.
+This table considers the multi-class task performed both directly and as a second step of the layered approach, without considering the definition of what a collective action problem is in the prompt.
+B.3 Additional Details on Topic Modeling
 Table B6 shows the complete configuration of topics identified in the set involving one thread from r/intj, one thread from r/wallstreetbets, and one thread from r/IronFrontUSA.
@@
 For completeness, we report the same visualization considering all topics in Figure B2.
-Table B4: List of topics identified by applying BERTopic on the reference dataset for topic modeling.
 Topic n. Topic keywords N. comments
@@
 3 robinhood, money, gme, app 105
-4 dick, twist, it, fuck 31
+4 d*ck, twist, it, f*ck 31
 5 this, fr, rarity, baaaa 25
@@
 8 government, white, targeted, person 10
+Table B6: List of topics identified by applying BERTopic on the reference dataset for topic modeling.
 Figure B2: Complete visual representation of the topic modeling application, resulting in 9 topics (plus noise).